# MetaToken: Detecting Hallucination in Image Descriptions by Meta Classification

### v1

@@ Abstract
Large Vision Language Models (LVLMs) have shown remarkable capabilities in multimodal tasks like visual question answering or image captioning. However, inconsistencies between the visual information and the generated text, a phenomenon referred to as hallucinations, remain an unsolved problem with regard to the trustworthiness of LVLMs. To address this problem, recent works proposed to incorporate computationally costly Large (Vision) Language Models in order to detect hallucinations on a sentence- or subsentence-level. In this work, we introduce MetaToken, a lightweight binary classifier to detect hallucinations on the token-level at negligible cost. Based on a statistical analysis, we reveal key factors of hallucinations in LVLMs which have been overseen in previous works. MetaToken can be applied to any open-source LVLM without any knowledge about ground truth data providing a reliable detection of hallucinations. We evaluate our method on four state-of-the-art LVLMs demonstrating the effectiveness of our approach.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) [4, 71, 51, 59, 7] gave rise to incorporate their few-shot learning capability into vision-language pre-trained models (VL-PTMs) [35, 33, 27, 50, 64, 24, 1] to learn superior cross-modal representations. These models [67, 17, 8, 73, 23], known as Large Vision Language Models (LVLMs), have demonstrated an impressive visual-language understanding by aligning text and visual features. However, besides their remarkable zero-shot performance on visual downstream tasks, LVLMs suffer from the problem of hallucinations [36, 40, 52] inherited from the underlying LLMs or caused by faulty interpretation of the image input by the vision branch.
@@
In the context of LVLMs, hallucination refers to the problem of inconsistencies between the generated text and the visual input [40], diminishing the trustworthiness of these models. Especially in safety-critical applications like autonomous driving [16, 57] or medicine [25, 32], the reliability of the underlying model is indispensable for decision making. In order to address this problem, recent works [15, 19, 70, 9, 66, 30, 72, 19, 62] have proposed additional instruction tuning datasets and pre-training strategies to detect and mitigate hallucinations on a sentence- or subsentence-level. Another common strategy comprises stacked L(V)LMs to post-hoc detect and rectify hallucinations [65, 68, 26, 6].
@@
Figure 1: MetaToken. Based on generated image captions (i), we calculate our proposed metrics (ii) (see Sec. 3.2). Afterwards, we apply the trained meta classifier (iii) to detect hallucinated and true objects (iv). Moreover, we show that (v) MetaToken can be easily integrated into LURE [72] to improve the hallucination mitigation.
@@
In this work, we tackle the problem of object hallucination in image captions. To this end, we introduce MetaToken, a lightweight hallucination detection method which can be applied to any open-source LVLM. MetaToken builds up on the idea of meta classification [38, 20, 5, 54, 53, 45, 13, 46, 55, 28] to detect hallucinated objects on the token-level based on the model output only. Fig. 1 depicts our approach. In contrast to existing methods, our approach neither requires an additional dataset, fine-tuning an L(V)LM nor cost-intensive L(V)LM prompting. Within a comprehensive statistical analysis, we investigate a broad set of metrics which are indicative of hallucinations providing deep insights into the sources of this specific type of model errors. Our examination reveals key indicators of hallucinations which have not been discovered in previous works. Moreover, our analysis shows that recently identified factors contain less information about hallucinations than our proposed metrics. We evaluate our method on four state-of-the-art LVLMs [8, 67, 73, 23] achieving area under receiver operator characteristic curve values [10] of up to $92.19\%$ and area under precision recall curve values [10] of up to $82.69\%$. Moreover, we show that our method can be incorporated into the LURE mitigation method [72]. While LURE reduces hallucinations by up to $52.98\%$, we achieve a hallucination reduction by up to $56.62\%$ through the superior precision-recall-ratio of MetaToken.
To sum up, our main contributions are as follows:
@@
  • We propose and investigate a comprehensive set of metrics as potential factors of object hallucinations.
  • Based on a statistical analysis, we demonstrate the insufficiency of recently proposed metrics and reveal key factors of object hallucinations.
@@
  • Based on these metrics, we introduce MetaToken, a lightweight binary classifier to reliably detect object hallucinations as a post-hoc method. MetaToken can be applied to any open-source LVLM without any knowledge about the ground truth data.
@@ 2 Related Work
2.1 Large Vision Language Models
VL-PTMs [35, 33, 27, 50, 64, 24, 1] have demonstrated remarkable performance in downstream task by learning universal cross-modal representations [11]. Recent advances in LLMs [4, 71, 51, 59, 7] gave rise to incorporate the few-shot learning capability of LLMs into VL-PTMs. These models [67, 17, 8, 73, 23], known as LVLMs, show an impressive visual-language understanding by aligning visual features with LLMs.
In general, LVLMs can be decomposed into three key components: (i) a vision encoder, which transforms the input image into vision tokens, (ii) a cross-modal alignment network, which aligns the visual and language information and (iii) an LLM, which generates the language output. The training of LVLMs usually consists of (i) the pre-training step using aligned vision-language pairs followed by (ii) vision-language instruction tuning enhancing the zero-shot performance on visual downstream tasks [31]. However, besides their notable success in vision-language tasks, LVLMs suffer from the problem of hallucinations [36] inherited from the underlying LLMs.
Table 1: Related Work on Hallucination Detection. A comparison of existing approaches on hallucination detection with respect to computational efficiency, i.e., whether the respective method requires an additional dataset, fine-tuning an L(V)LM or prompting an L(V)LM, and the ability of the respective method to detect hallucinations on different semantic levels (hallucination type). indicates ’yes’, indicates ’no’.
computational efficiency hallucination type
method w/o add. dataset w/o fine-tuning w/o prompting object attribute sentence
LogicCheckGPT [65]
Woodpecker [68]
M-HalDetect [19]
HaELM [62]
FAITHSCORE [26]
UNIHD [6]
Ours
2.2 Object Hallucination
@@
The phenomenon of object hallucination refers to the problem of inconsistencies between the generated text and the visual input [40]. Generally speaking, hallucinations in LVLMs can occur on different semantic levels, where coarse-grained object hallucination [52] refers to objects generated in the language output, which are not depicted in the input image, whereas fine-grained hallucination describes inconsistencies with respect to object attributes or relations between objects [36, 40]. For a comprehensive survey on hallucinations in LVLMs, we refer to [40].
The problem of hallucination mitigation is mainly tackled by either retraining the model with an instruction tuning dataset [15, 19], rectifying image captions as a post-processing step or incorporating new pre-training or generation strategies. LURE [72] serves as a post-hoc method to rectify object hallucinations by training an LVLM-based revisor to reconstruct less hallucinatory descriptions. MARINE [70] enriches the visual context of LVLMs by incorporating object grounding features into the LLM input. In [9], a new pre-training objective is introduced to mitigate object hallucinations by improving the object-level image-text alignment. Motivated by the idea of unlearning, EFUF [66] employs gradient ascent to penalize hallucinations during the unlearning process. Moreover, Visual Contrastive Decoding [30] leverages distorted visual inputs to contrast the output distributions of original and distorted images.
@@
Since the reliability of the generated language output still remains an unsolved problem, many researches focus on the problem of hallucination detection. In [65], the problem of hallucination detection and mitigation is solved simultaneously by raising logical correlated questions and checking for logical consistency throughout the generated answers afterwards. Similarly, Woodpecker [68] serves as a post-processing hallucination detection and correction method incorporating visual knowledge validation for both instance- and attribute-level hallucinations. A human-labeled dataset is published in [19], which is used to train a binary classifier based on the LVLM InstructBLIP [8] to classify between accurate and inaccurate sentences. In [62], an LLM-based hallucination evaluation framework is introduced by training an LLM to distinguish between hallucinated and hallucination-free image captions. Both, [26] and [6], propose a pipeline consisting of several LVLMs and LLMs to verify each claim contained in the generated language output. In this work, we tackle the problem of object hallucination detection by training a lightweight binary classifier. Our method serves as a post-hoc method to classify between true and hallucinated objects. In contrast to existing methods, our approach neither requires an additional dataset, fine-tuning an L(V)LM nor cost-intensive L(V)LM prompting for claim verification (see Tab. 1).
@@ 2.3 Meta Classification
In classical machine learning, meta classification refers to the problem of how to best combine predictions from an ensemble of classifiers [38]. In terms of deep learning, this concept has been transferred to the classification whether a prediction is true or false based on uncertainty metrics [20]. Several works have applied this idea to natural language processing [60, 41, 18], image classification [5], semantic segmentation [54, 53, 45, 13], video instance segmentation [46] and object detection [55, 28]. We are the first to transfer the idea of meta classification to the problem of hallucination detection for LVLMs. Based on a statistical analysis of key factors of object hallucination in LVLMs, we identify input metrics outperforming classical uncertainty-based metrics.
@@ 2.4 Hallucination Evaluation
Since different studies [52, 9] have shown that standard image captioning metrics like BLEU [49], METEOR [29], CIDEr [61] and SPICE [2] are not capable of measuring object hallucinations properly, most works on hallucination mitigation measure the performance of their proposed method in terms of the Caption Hallucination Assessment with Image Relevance (CHAIR) metric [52]. The CHAIR metric measures the proportion of hallucinated MSCOCO objects [37] in an image caption by matching the MSCOCO objects in the generated text against the ground truth objects provided in the MSCOCO image captioning and object detection datasets [37]. Thus, for every generated MSCOCO object, the CHAIR metric provides a binary label indicating whether the object is true, i.e., contained in the image, or hallucinated. Further datasets and evaluation methods have been proposed to evaluate the performance of LVLMs across multiple multi-modal tasks [42, 15, 14, 19, 43, 39, 63, 69, 62, 36, 21]. While some of the proposed evaluation methods ask LLMs to output quality-related scores [42, 15, 69] or measure the image-text similarity [21], other methods use a prompt template to query hallucination-related questions and force the model to answer either ’yes’ or ’no’. POPE [36] initially proposed to prompt LVLMs with the template "Is there a <obj> in the image?" to evaluate the model’s performance with respect to object hallucinations. In [14] and [63], this idea is extended by including questions on attribute-level. NOPE [43] comprises negative pronoun data only for different visual question answering tasks. However, the results in [36] and [14] have shown that LVLMs tend to answer ’yes’, which results in a low recall for hallucinated objects. Moreover, the LLM- and similarity-based scores [42, 15, 69, 21] evaluate the entire image caption in terms of continuous scores instead of providing a binary label for each generated object. Thus, in this work, we rely on the CHAIR metric to evaluate object hallucinations.
@@ 3 Method
The aim of our method is to detect object-level hallucinations [52] in the text output of LVLMs. To this end, we build metrics based on information retrieved from the model output that have been shown to correlate with object hallucinations. These metrics are used to train a lightweight binary meta model to classify between hallucinated and true objects. At inference time, we can detect hallucinations by computing the proposed metrics and applying the trained meta model afterwards.
@@ 3.1 Notation
We denote the sequence of generated output tokens by $s=(t_{0},\dots,t_{K})$ with sequence length $K+1$. Moreover, let $s_{i}=(t_{0},\dots,t_{i})$ denote the generated output sequence at generation step $i$. The probability of generating token $t_{i+1}\in\mathcal{V}$ at generation step $i+1$ given the input image $x$, the provided prompt $q$ and the already generated tokens $s_{i}$ can then be formulated as $p(t_{i+1}|x,q,s_{i})$. For a shorter notation, we define $\hat{p}_{i+1}=p(t_{i+1}|x,q,s_{i})$. Furthermore, let $p_{i+1}$ denote the probability distribution at generation step $i+1$ over the dictionary $\mathcal{V}$ and $|\mathcal{V}|$ the cardinality of $\mathcal{V}$.
Given the language output $s$, we can extract all MSCOCO objects [37] contained in the generated text using the CHAIR method [52]. Let $\mathcal{C}=\{c_{0},\dots,c_{79}\}$ denote the set of all $80$ MSCOCO objects. We denote the set of all MSCOCO objects contained in the sequence $s$ by $\mathcal{O}_{s}=\{o_{0},\dots,o_{z}\}$ with $o_{j}\in\mathcal{C}$ for all $j=0,\dots,z$. Since the generated string of an object might consist of several tokens, we define for every object $o_{j}\in\mathcal{O}_{s}$ the start token $t_{o_{j,s}}$ at position $0\leq o_{j,s}\leq K$ as well as the end token $t_{o_{j,e}}$ at position $0\leq o_{j,e}\leq K$. Finally, we denote the set of all captions which mention object $o_{j}$ as $\mathcal{S}_{o_{j}}$ with cardinality $|\mathcal{S}_{o_{j}}|$.
3.2 Input Metrics
Recent works [52, 36, 62, 72] have investigated influencing factors of object hallucinations. First, the analysis in [36] has shown that LVLMs are prone to hallucinate objects from the underlying visual instruction tuning datasets. Since the MSCOCO dataset [37] is widely used for instruction tuning, LVLMs tend to hallucinate frequent MSCOCO objects. Furthermore, LVLMs have a high hallucination rate on co-occurring objects, that is, objects which co-occur in the visual instruction tuning datasets frequently tend to occur in the generated language output of LVLMs together even though only one of the objects exists in the image. Thus, we consider the MSCOCO class index (Eq. 1) which maps a list of synonyms [44] to one of the $80$ MSCOCO classes as well as the co-occurrence score (Eq. 2) inspired from [72]. Second, the results in [62] indicate that LVLMs often generate true segments at the beginning while the risk of hallucinations increases at the letter part of the generated responses. Thus, we take account of the relative position (Eq. 3) of a generated object and the absolute occurrence (Eq. 4) of the object in the generated text. Third, to account for the over-reliance of LVLMs on language priors during the generation process [52, 62], we consider the mean absolute attention on the image tokens (Eq. 5). Finally, we regard the model uncertainty through different dispersion measures (Eq. 7-(13)) which have been shown to correlate with model errors in different fields [53, 55, 60]. For a generated object $o_{j}\in\mathcal{O}_{s}$ from the output sequence $s=(t_{0},\dots,t_{K})$, we define
  • the MSCOCO class index
    $I_{o_{j}}=k,\quad\textrm{such that }o_{j}=c_{k},$ (1)
  • the co-occurrence score
    $H_{o_{j}}=\sum_{\begin{subarray}{c}l=0\\ o_{l}\neq o_{j}\end{subarray}}^{z}\frac{|\mathcal{S}_{o_{l}}\cap\mathcal{S}_{o_{j}}|}{|\mathcal{S}_{o_{l}}\cup\mathcal{S}_{o_{j}}|},$ (2)

### v2

@@ Abstract
Large Vision Language Models (LVLMs) have shown remarkable capabilities in multimodal tasks like visual question answering or image captioning. However, inconsistencies between the visual information and the generated text, a phenomenon referred to as hallucinations, remain an unsolved problem with regard to the trustworthiness of LVLMs. To address this problem, recent works proposed to incorporate computationally costly Large (Vision) Language Models in order to detect hallucinations on a sentence- or subsentence-level. In this work, we introduce MetaToken, a lightweight binary classifier to detect hallucinations on the token-level at negligible cost. Based on a statistical analysis, we reveal key factors of hallucinations in LVLMs. MetaToken can be applied to any open-source LVLM without any knowledge about ground truth data providing a calibrated detection of hallucinations. We evaluate our method on four state-of-the-art LVLMs demonstrating the effectiveness of our approach.
@@ 1 INTRODUCTION
LVLMs have demonstrated an impressive visual-language understanding by aligning text and visual features. However, besides their remarkable zero-shot performance on visual downstream tasks, LVLMs suffer from the problem of hallucinations [Li et al., 2023b, Liu et al., 2024b, Rohrbach et al., 2018] inherited from the underlying LLMs or caused by faulty interpretation of the image input by the vision branch. In the context of LVLMs, hallucination refers to the problem of inconsistencies between the generated text and the visual input [Liu et al., 2024b], diminishing the trustworthiness of these models. Especially in safety-critical applications like autonomous driving [Gao et al., 2024, Tian et al., 2024] or medicine [Jiang et al., 2024, Li et al., 2023a], the reliability of the underlying model is indispensable for decision making. In order to address this problem, recent works [Liu et al., 2024a, Gunjal et al., 2023, Zhao et al., 2024, Dai et al., 2023b, Xing et al., 2024] have proposed additional instruction tuning datasets and pre-training strategies to detect and mitigate hallucinations on a sentence- or subsentence-level. Another common strategy comprises stacked L(V)LMs to post-hoc detect and rectify hallucinations [Wu et al., 2024, Yin et al., 2023, Jing et al., 2023].
@@
Figure 1: MetaToken. Based on generated image captions (i), we calculate our proposed input features (ii) (see Sec. 3.2). Afterwards, we apply the trained meta classifier (iii) to detect hallucinated and true objects (iv). Moreover, (v) MetaToken can be easily integrated into LURE [Zhou et al., 2023] to improve the hallucination mitigation.
Table 1: Related Work on Hallucination Detection. A comparison of existing approaches on hallucination detection with respect to computational efficiency, i.e., whether the respective method can be implemented without an additional dataset, fine-tuning or prompting an L(V)LM. indicates ’yes’, indicates ’no’.
method w/o add. dataset w/o fine-tuning w/o prompting
LogicCheckGPT [Wu et al., 2024]
Woodpecker [Yin et al., 2023]
M-HalDetect [Gunjal et al., 2023]
HaELM [Wang et al., 2023]
FAITHSCORE [Jing et al., 2023]
UNIHD [Chen et al., 2024]
Ours
@@
In this work, we tackle the problem of object hallucination in image captions. To this end, we introduce MetaToken, a lightweight hallucination detection method which can be applied to any open-source LVLM. MetaToken builds up on the idea of meta classification [Lin and Hauptmann, 2003, Hendrycks and Gimpel, 2017, Chen et al., 2019, Rottmann et al., 2020, Fieback et al., 2023] to detect hallucinated objects on the token-level based on the model output only. Fig. 1 depicts our approach. In contrast to existing methods, our approach neither requires an additional dataset, fine-tuning an L(V)LM nor cost-intensive L(V)LM prompting. Within a comprehensive statistical analysis, we investigate a broad set of input features which are indicative of hallucinations providing deep insights into the sources of this specific type of model errors. We evaluate our method on four state-of-the-art (SOTA) LVLMs [Dai et al., 2023a, Ye et al., 2023, Zhu et al., 2023, Huang et al., 2023] achieving area under receiver operator characteristic curve values [Davis and Goadrich, 2006] of up to $92.12\%$ and area under precision recall curve values [Davis and Goadrich, 2006] of up to $84.01\%$. Moreover, we show that our method can be incorporated into the LURE mitigation method [Zhou et al., 2023]. While LURE reduces hallucinations by up to $52.98\%$, we achieve a hallucination reduction by up to $56.62\%$ through the superior precision-recall-ratio of MetaToken. Our main contributions are as follows:
@@
  • We propose and investigate a comprehensive set of statistics as potential factors of object hallucinations.
@@
  • Based on these statistics, we introduce MetaToken, a lightweight binary classifier to detect object hallucinations as a post-hoc method. MetaToken can be applied to any open-source LVLM without any knowledge about the ground truth data.
@@ 2 RELATED WORK
2.1 Hallucinations in LVLMs
@@
Hallucinations in LVLMs can occur on different semantic levels, where coarse-grained object hallucination [Rohrbach et al., 2018] refers to objects generated in the language output, which are not depicted in the input image, whereas fine-grained hallucination describes inconsistencies with respect to object attributes or relations between objects [Li et al., 2023b, Liu et al., 2024b]. For a comprehensive survey on hallucinations in LVLMs, we refer to [Liu et al., 2024b].
@@
The problem of hallucination mitigation is mainly tackled by either retraining the model with an instruction tuning dataset [Liu et al., 2024a, Gunjal et al., 2023], rectifying image captions as a post-processing step or incorporating new pre-training or generation strategies. LURE [Zhou et al., 2023] serves as a post-hoc method to rectify object hallucinations by training an LVLM-based revisor to reconstruct less hallucinatory descriptions. MARINE [Zhao et al., 2024] enriches the visual context of LVLMs by incorporating object grounding features into the LLM input. In [Dai et al., 2023b], a new pre-training objective is introduced to mitigate object hallucinations by improving the object-level image-text alignment. Since the reliability of the generated language output still remains an unsolved problem, many researches focus on the problem of hallucination detection. In [Wu et al., 2024], the problem of hallucination detection and mitigation is solved simultaneously by raising logical correlated questions and checking for logical consistency throughout the generated answers afterwards. Similarly, Woodpecker [Yin et al., 2023] serves as a post-processing hallucination detection and correction method incorporating visual knowledge validation for both instance- and attribute-level hallucinations. A human-labeled dataset is published in [Gunjal et al., 2023], which is used to train a binary classifier based on the LVLM InstructBLIP [Dai et al., 2023a] to classify between accurate and inaccurate sentences. In [Wang et al., 2023], an LLM-based hallucination evaluation framework is introduced by training an LLM to distinguish between hallucinated and hallucination-free image captions. Both, [Jing et al., 2023] and [Chen et al., 2024], propose a pipeline consisting of several LVLMs and LLMs to verify each claim contained in the generated language output. In this work, we tackle the problem of object hallucination detection using meta classification, that is, training a lightweight binary classifier to distinguish between true and hallucinated objects. In contrast to existing methods, our approach neither requires an additional dataset, fine-tuning an L(V)LM nor cost-intensive L(V)LM prompting for claim verification (see Tab. 1).
@@ 2.2 Meta Classification
In classical machine learning, meta classification refers to the problem of how to best combine predictions from an ensemble of classifiers [Lin and Hauptmann, 2003]. In terms of deep learning, this concept has been transferred to the classification whether a prediction is true or false based on uncertainty features [Hendrycks and Gimpel, 2017]. Several works have applied this idea to natural language processing [Vasudevan et al., 2019, Liu et al., 2022, Gui et al., 2024], image classification [Chen et al., 2019], semantic segmentation [Rottmann et al., 2020, Rottmann and Schubert, 2019, Maag et al., 2020, Fieback et al., 2023], video instance segmentation [Maag et al., 2021] and object detection [Schubert et al., 2021, Kowol et al., 2020]. We are the first to transfer the idea of meta classification to the problem of hallucination detection for LVLMs. Based on a statistical analysis of key factors of hallucinations in LVLMs, we identify input features outperforming classical uncertainty-based statistics.
@@ 2.3 Hallucination Evaluation
Since different studies [Rohrbach et al., 2018, Dai et al., 2023b] have shown that standard image captioning metrics like BLEU [Papineni et al., 2002], METEOR [Lavie and Agarwal, 2007], CIDEr [Vedantam et al., 2015] and SPICE [Anderson et al., 2016] are not capable of measuring hallucinations properly, most works on hallucination mitigation measure the performance of their proposed method in terms of the Caption Hallucination Assessment with Image Relevance (CHAIR) metric [Rohrbach et al., 2018]. The CHAIR metric measures the proportion of hallucinated objects in an image caption by matching the objects in the generated text against the ground truth annotations [Lin et al., 2014]. Thus, for every generated object, the CHAIR metric provides a binary label indicating whether the object is true, i.e., contained in the image, or hallucinated. Further datasets and evaluation methods have been proposed to evaluate the performance of LVLMs across multiple multi-modal tasks [Fu et al., 2023, Gunjal et al., 2023, Lovenia et al., 2024, Liu et al., 2023]. While some of the proposed evaluation methods ask LLMs to output quality-related scores [Liu et al., 2025, Liu et al., 2024a, Yu et al., 2024] or measure the image-text similarity [Hessel et al., 2021], other methods [Li et al., 2023b, Fu et al., 2023, Wang et al., 2024] use a prompt template to query hallucination-related questions and force the model to answer either ’yes’ or ’no’. However, the results in [Li et al., 2023b] and [Fu et al., 2023] have shown that LVLMs tend to answer ’yes’, which results in a low recall for hallucinated objects. Moreover, the LLM- and similarity-based scores [Liu et al., 2025, Liu et al., 2024a, Yu et al., 2024, Hessel et al., 2021] evaluate the entire image caption in terms of continuous scores instead of providing a binary label for each generated object. Thus, we rely on the CHAIR metric to evaluate hallucinations.
@@ 3 METHOD
The aim of our method is to detect hallucinations in the text output of LVLMs leveraging the idea of meta classification. To this end, we build input features based on the model output that have been shown to correlate with hallucinations. These features are used to train a lightweight binary meta model to classify between hallucinated and true objects. At inference time, we can detect hallucinations by computing the proposed features and applying the trained meta model afterwards. A formal definition of meta classification is provided in Sec. 3.3.
@@ 3.1 Notation
We denote the sequence of generated output tokens by $s=(t_{0},\dots,t_{K})$ with sequence length $K+1$. Moreover, let $s_{i}=(t_{0},\dots,t_{i})$ denote the generated output sequence at generation step $i$. The probability of generating token $t_{i+1}\in\mathcal{V}$ at generation step $i+1$ given the input image $x$, the provided prompt $q$ and the already generated tokens $s_{i}$ can be formulated as $p(t_{i+1}|x,q,s_{i})$. For a shorter notation, we define $\hat{p}_{i+1}=p(t_{i+1}|x,q,s_{i})$. Furthermore, let $p_{i+1}$ denote the probability distribution at generation step $i+1$ over the dictionary $\mathcal{V}$ and $|\mathcal{V}|$ the cardinality of $\mathcal{V}$.
Given the language output $s$, we extract all objects contained in the generated text. We denote the set of objects contained in the sequence $s$ by $\mathcal{O}_{s}=\{o_{0},\dots,o_{z}\}$. Since the generated string of an object might consist of several tokens, we define for every object $o_{j}\in\mathcal{O}_{s}$ the start token $t_{o_{j,s}}$ at position $0\leq o_{j,s}\leq K$ as well as the end token $t_{o_{j,e}}$ at position $0\leq o_{j,e}\leq K$.
3.2 Input Features
Recent works [Rohrbach et al., 2018, Wang et al., 2023] have investigated influencing factors of object hallucinations. First, the results in [Wang et al., 2023] indicate that LVLMs often generate true segments at the beginning while the risk of hallucinations increases at the letter part of the generated responses. Thus, we take account of the relative position (Eq. 1) of a generated object and the absolute occurrence (Eq. 2) of the object in the generated text. Second, to account for the over-reliance of LVLMs on language priors during the generation process [Rohrbach et al., 2018, Wang et al., 2023], we consider the mean absolute attention on the image tokens (Eq. 3). Finally, we regard the model uncertainty through different dispersion measures (Eq. 4-(11)) which have been shown to correlate with model errors in different fields [Rottmann and Schubert, 2019, Schubert et al., 2021, Vasudevan et al., 2019] including the sequence score [^1] https://huggingface.co/docs/transformers/main/en/main˙classes/text˙generation#transformers.GenerationMixin.compute˙transition˙scores calculated during the LVLM generation process. For a generated object $o_{j}\in\mathcal{O}_{s}$ from the output sequence $s=(t_{0},\dots,t_{K})$, we define